Most of the galaxies in the universe are observed to be moving away from Earth. Suppose a particular galaxy emits orange light with a frequency of 5.000×1014Hz.

If the galaxy is receding from Earth with a speed of 4875 km/s, what is the frequency of the light when it reaches Earth?
Answer #1

As we know that;

= f= 5*10^14 Hz

= v= 4874 km/s = 4.875*10^6 m/s

= c= 3*10^8 m/s

Then,